# Energy, enstrophy and helicity transfers in polymeric turbulence

### v1

@@ 4 Scale energy
In this section, we detail the influence of the polymeric additives on the scale-by-scale energy transfer. We first focus on the budget equation for $\left\langle{\delta q^{2}}\right\rangle$ and look at the modifications of the average energy cascade. This is the $r$-space equivalent of the Fourier analysis by Rosti et al. (2023). Then, we look in detail at the nonlinear ($\Phi_{f}^{\delta q^{2}}$) and polymeric ($\Phi_{p}^{\delta q^{2}}$) fluxes to highlight how polymers modify the intermittent nature of the energy transfers.
4.1 Scale-by-scale budget
@@
Figure 5: Scale-by-scale energy budget. The panels are for $De=0$, $De=1/9$, $De=1/3$, $De=1$, $De=3$ and $De=9$, from top-left to bottom-right. The yellow and blue shaded regions identify the inertial and elastic ranges of scales.
@@
Figure 5 plots the scale-by-scale energy budget for both HIT and PHIT. We start looking at the HIT case ($De=0$), which depicts the Richardson-Kolmogorov picture of turbulence (Richardson, 1922). Energy is injected at the largest scales $r>r_{e}\approx 1.84$ by the external forcing $\text{d}\langle F^{\delta q^{2}}\rangle/\text{d}r\approx(4/3)\langle\varepsilon_{f}^{\delta q^{2}}\rangle$, and is then transferred from larger to smaller scales within the inertial range $0.043\approx r_{d}\leq r\leq r_{e}$ via the classical cascade process; $r_{e}$ denotes the scale above which the forcing term dominates (i.e. separates the energetic and inertial ranges of scales), while $r_{d}$ denotes the scale below which the dissipative term dominates (i.e. separates the inertial and dissipative ranges of scales). As predicted by the Kármán-Howarth equation (de Kármán & Howarth, 1938), the energy transfer rate is an invariant in the inertial range of scales and equals the dissipation rate $\text{d}\langle\Phi_{f}^{\delta q^{2}}\rangle/\text{d}r\approx-(4/3)\langle\varepsilon_{f}^{\delta q^{2}}\rangle$; see the pleateau in the top left panel in figure 5. At the smallest scales $r<r_{d}$ the viscous contribution $D_{v}^{\delta q^{2}}$ dominates, and $\left\langle{\delta q^{2}}\right\rangle$ is dissipated into heat by viscous friction.
@@
Polymeric additives introduce an alternative route for energy transfer among scales ($\Phi_{p}^{\delta q^{2}}$) and an additional dissipative process ($D_{p}^{\delta q^{2}}$). For small $De$($<1$), the global picture resembles what observed in HIT. In this case, indeed, the polymer chains are not effectively stretched as their relaxation time is small compared to the characteristic time scales of the flow, and they quickly relax back to their equilibrium lengths; here, $\Phi_{p}^{\delta q^{2}}\approx D_{p}^{\delta q^{2}}\approx\pi^{\delta q^{2}}\approx 0$. As $De$ increases ($De\geq 1$) the relevance of the non-Newtonian contribution increases, and the polymeric additives modify the energy cascade in a non trivial way. Energy enters the system at the large scales ($r>r_{e}$) and is dissipated away at the same rate, via both the fluid and polymeric dissipation, i.e.
$\frac{\text{d}}{\text{d}r}\left\langle{F^{\delta q^{2}}}\right\rangle\approx\frac{4}{3}\left\langle{\varepsilon_{f}^{\delta q^{2}}+\pi^{\delta q^{2}}}\right\rangle.$
@@
In the intermediate range of scales ($r_{d}<r<r_{e}$) energy is then transferred from larger to smaller scales by two alternative routes, respectively associated with (i) the classic inertia-driven energy cascade $\Phi_{f}^{\delta q^{2}}$ and (ii) the fluid/polymer interaction $\Phi_{p}^{\delta q^{2}}$. Here the viscous effects are negligible and the total energy transfer rate is an invariant and matches the total dissipation rate, i.e.
$\frac{\text{d}}{\text{d}r}\left(\left\langle{\Phi_{p}^{\delta q^{2}}}\right\rangle+\left\langle{\Phi_{f}^{\delta q^{2}}}\right\rangle\right)\sim-\frac{4}{3}\left\langle{\varepsilon_{f}^{\delta q^{2}}+\pi^{\delta q^{2}}}\right\rangle.$
Eventually, at the smallest scales ($r<r_{d}$) energy is dissipated away by both the fluid and the polymers, and
$\frac{\text{d}}{\text{d}r}\left\langle{D_{f}^{\delta q^{2}}}\right\rangle\rightarrow\frac{4}{3}\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle\ \text{and}\ \frac{\text{d}}{\text{d}r}\left\langle{D_{p}^{\delta q^{2}}}\right\rangle\rightarrow\frac{4}{3}\left\langle{\pi^{\delta q^{2}}}\right\rangle\ \text{for}\ r\rightarrow 0.$
@@
The range of scales where the energy cascade is effectively modified by the presence of the polymers changes with $De$; see the blue shaded region in figure 5. One can identify for each $De$ the scale $r_{p}^{*}$ below which elastic effects are expected to dominate using simple arguments. We estimate this scale by comparing the local turnover time of eddies of size $r$, i.e. $\tau_{f}(r)\sim r/\left\langle{\delta V^{2}}\right\rangle^{1/2}$ — where $\delta V(r)=(u_{i}(\bm{x}+\bm{r})-u_{i}(\bm{x}))\cdot r_{i}/r)$ — with the polymeric relaxation time $\tau_{p}$. Thus, $r_{p}^{*}$ is estimated as:
@@
For $r>r_{p}^{*}$ the time scale of the fluid fluctuations is larger than $\tau_{p}$: the polymers are only marginally stretched and thus their influence on the flow is weak. On the contrary, when $r<r_{p}^{*}$ the time scale of the fluid is smaller than the polymeric relaxation time: the polymers are stretched more effectively at these scales and thus modify the organisation of the corresponding fluctuations. For the present cases, we measure $r_{p}^{*}\approx 0.04,0.32,1.34$ and $2.81$ for $De=1/9,1/3,1$ and $3$: the range of scales where elastic effects significantly modify the energy cascade increases with $De$, in agreement with the widening of the elastic range of scales shown in figure 5. For $De=9$, $\tau_{p}>r/\left\langle{\delta V^{2}}\right\rangle^{1/2}$ for all $r$, suggesting that the polymeric additives effectively modify the energy cascade in the whole range of scales. We reiterate that when $De$ is very large and $\tau_{p}\gg r/\left\langle{\delta V^{2}}\right\rangle^{1/2}$ the polymeric chains decouple from the carrier phase and their influence on the carrier flow is rather low.
The multiple scaling behaviour observed in figure 1 for $De\geq 1$ is consistent with $r_{p}^{*}$ being well within the $r_{e}<r<r_{d}$ range of scales (where both the viscous effects and the external forcing are negligible) for large enough Deborah and Reynolds numbers. The energy cascade driven by the fluid inertia $\Phi_{f}^{\delta q^{2}}$ dominates at the larger scales $r_{p}^{*}\lessapprox r\lessapprox r_{e}$, where indeed $\left\langle{\delta q^{2}}\right\rangle\sim r^{2/3}$. The cascade driven by the fluid/polymer interaction $\Phi_{p}^{\delta q^{2}}$, instead, dominates at the smaller $r_{d}\lessapprox r\lessapprox r_{p}^{*}$ scales, where $\left\langle{\delta q^{2}}\right\rangle\sim r^{\xi}$ with $\xi\approx 1.3\approx 4/3$. Accordingly, this peculiar multiscaling behaviour is not observed for $De<1$ and/or small $Re$, as in these cases $r_{p}^{*}$ is within the dissipative range of scales (see for example De Angelis et al., 2005; Perlekar et al., 2010).
A last comment regards the influence of $De$ on the dissipative range of scales. Our data show that $r_{d}$ (i.e. scale below which the total dissipative term $D_{v}^{\delta q^{2}}+D_{p}^{\delta q^{2}}$ dominates) decreases with $De$, thus indicating that the dissipative range of scales progressively shrinks as the polymeric relaxation time increases.
4.2 The nonlinear and non Newtonian fluxes
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ 4.2 The nonlinear and non Newtonian fluxes
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 7: Dependence of the distribution of the non Newtonian flux $\Phi_{p}$ on the Deborah number, for (left) $r=0.1$, (centre) $r=0.3$, and (right) $r=0.9$. Top: distribution of $\Phi_{p}^{\delta q^{2}}$. Bottom: distribution of $\Phi_{p}^{\delta q^{2}}/\Phi_{p,rms}^{\delta q^{2}}$. In the bottom panels the inset provide a zoom of the normalised distributions.
@@
The average picture of the cascade and the interscale exchanges in the $r-$space is however not representative of the actual physical processes. In the previous section, we have shown the influence of the polymers on the energy cascade in an average sense only, even though it is known that the interscale energy transfer is highly intermittent (see for example Piomelli et al., 1991; Domaradzki et al., 1993; Cerutti & Meneveau, 1998). Moreover, even though energy cascade in HIT is on an average from larger to smaller scales, there exist localised regions in the flow where energy actually cascades from smaller to larger scales, opposite to the average sense. We now investigate how the presence of polymeric additives modifies this picture, and whether this localised inverse energy transfer is also detected in the transfer mechanism driven by the fluid/polymer interaction.
@@
Figures 6 and 7 plot the probability distribution functions of $\Phi_{f}^{\delta q^{2}}$ and $\Phi_{p}^{\delta q^{2}}$ for different $De$. We consider three different separations $r$, i.e. $r=0.1$ where $\Phi_{p}^{\delta q^{2}}$ dominates for $De>1$, and $r=0.3,0.9$ where $\Phi_{f}^{\delta q^{2}}$ dominates. In the present convention, a positive $\Phi^{\delta q^{2}}>0$ indicates backscatter, i.e. local events where energy goes from smaller to larger scales, while a negative $\Phi^{\delta q^{2}}<0$ indicates events where energy goes from larger to smaller scales. The distributions reveal the highly intermittent and non-Gaussian nature of both $\Phi_{f}^{\delta q^{2}}$ and $\Phi_{p}^{\delta q^{2}}$, given their heavy tails (see Ishihara et al., 2009; Yasuda & Vassilicos, 2018, for HIT): backscatter and extreme forward events exist with much higher probability than for a normal distribution. The asymmetry of the distributions shows that on average both $\Phi_{f}^{\delta q^{2}}$ and $\Phi_{p}^{\delta q^{2}}$ are negative, and that the average interscale transfer of energy is in the forward sense for both routes. To be quantitative, for $De=1$ and $r=0.9$ ($r=0.1$) the probability of $\Phi_{f}^{\delta q^{2}}>\Phi^{\delta q^{2}}_{f,rms}$ is $0.0672$ ($0.0274$), while the probability of events with $\Phi_{f}^{\delta q^{2}}<-\Phi_{f,rms}^{\delta q^{2}}$ is $0.1640$ ($0.0767$); similarly, the probability of $\Phi_{p}^{\delta q^{2}}>\Phi_{p,rms}^{\delta q^{2}}$ is $0.0323$ ($0.0105$), while the probability of events with $\Phi_{p}^{\delta q^{2}}<-\Phi_{p,rms}^{\delta q^{2}}$ is $0.2402$ ($0.1545$).
We now focus on how polymers influence the tails of the distributions, i.e. the extreme events. We start looking at the nonlinear flux $\Phi_{f}^{\delta q^{2}}$ (see figure 6). For the three chosen scales, the addition of polymers leads to narrower tails: extreme events of both direct and inverse energy transfers along the classical route are inhibited. This effect becomes more evident as $De$ increases, in agreement with a smoother velocity field, and with the larger amount of energy that is redirected towards the polymer-driven transfer route. When looking at the fluxes normalised with their root-mean-square value (bottom panels), we note that the distributions at $r=0.3$ and $r=0.9$ collapse reasonably well for all $De$. This means that in the $\Phi_{f}^{\delta q^{2}}$ dominated range, the presence of the polymers modifies the strength of the extreme cascade events, but does not influence its intermittency. In other words, in the range of scales where the polymeric contribution is weak the self-similarity of the classical cascade process is preserved. At the small scales $r=0.1$ where $\Phi_{p}^{\delta q^{2}}$ dominates, however, the normalised distributions of $\Phi_{f}^{\delta q^{2}}$ do not overlap: here polymers modify the intermittency of the inertia-driven cascade. The nonlinear flux $\Phi_{f}^{\delta q^{2}}$ has a small average contribution at this scale, whose intermittency increases with $De$. For very large elasticity, the tails begin to shrink again as polymers decouple from these small scales easily thus reducing their effect on small-scale fluctuations.
Figure 7 considers the distribution of the polymeric flux $\Phi_{p}^{\delta q^{2}}$ at different scales. An increase of $De$ results in wider tails, with the asymmetry of the distribution increasing with $De$ in agreement with an increase of the average value of $\Phi_{p}^{\delta q^{2}}$. When considering the normalised flux (bottom panels), the scenario changes compared to $\Phi_{f}^{\delta q^{2}}$. In this case, indeed, the distributions overlap reasonably well at the small scale $r=0.1$, but not for the larger $r=0.3$ and $r=0.9$ scales. An increase of $De$ modifies the space-time intermittency of the interscale energy transfer driven by the polymeric microstructure at large scales only, while self-similarity with $De$ is observed at small $r$ in the $\Phi_{p}^{\delta q^{2}}$ dominated range. Accordingly with the increased asymmetry, the larger $De$ corresponds to a narrower positive tails and to a slightly heavier negative tail.
Figure 8: Distribution of $\cos(\theta)=\bm{u}^{*}\cdot(\delta\bm{\omega}\times\delta\bm{u})/(||\bm{u}^{*}||||\delta\bm{u}\times\delta\bm{\omega}||)$ for different scales. In order from top-left to bottom-right, the panels are for $De=0$, $De=1/9$, $De=1/3$, $De=1$, $De=3$ and $De=9$.
@@
We now provide a further insight on the influence of the polymers on the inertia-driven energy cascade process by looking at the simple relation introduced by Baj et al. (2022) that links the nonlinear flux with the Lamb vector $\delta\bm{\omega}\times\delta\bm{u}$. Exploiting the Lamb decomposition and manipulating the equations, it is indeed possible to show that
@@
Since $||\delta\bm{u}\times\delta\bm{\omega}||^{2}=||\delta\bm{u}||^{2}||\delta\bm{\omega}||^{2}-||\delta h||^{2}$, it is clear that there is an implicit connection between the scale-by-scale helicity $\delta h$ and the interscale energy transfer. Indeed, a large $\delta h=\delta\bm{\omega}\cdot\delta\bm{u}$ corresponds to a small $\delta\bm{\omega}\times\delta\bm{u}$, and therefore to a weaker local transfer of energy. This agrees with the observation of previous authors which found that the magnitude of the helicity has an impact on the local transfer of energy (see for example Pelz et al., 1985; Stepanov et al., 2015). To gain further insights on how polymers influence the interscale energy transfer, we look at the angle $\theta$ between $\bm{u}^{*}$ and $\delta\bm{\omega}\times\delta\bm{u}$ such that
$\cos(\theta)=\frac{\bm{u}^{*}\cdot\left(\delta\bm{\omega}\times\delta\bm{u}\right)}{||\bm{u}^{*}||||\delta\bm{u}\times\delta\bm{\omega}||}.$ (22)
Figure 8 plots the distributions of $\cos(\theta)$ at different scales $r$ for $0\leq De\leq 9$. We start our discussion with HIT, where the distribution is positively skewed for all $r$. This is consistent with equation 21, which implies that on average the left hand side ($||\bm{u}^{*}||||\delta\bm{\omega}\times\delta\bm{u}||\cos(\theta)$) is positive. Similar to what found by Baj et al. (2022), as the separation increases the probability of parallel ($\cos(\theta)=1$) and antiparallel ($\cos(\theta)=-1$) alignment increases, with a larger probability of the parallel events. For PHIT ($De>0$), the picture changes in a scale-dependent manner. At small $De(\ll 1)$, as expected, the distributions remain similar to HIT. For large $De(\gtrsim 1)$, instead, at small scales $r$ where the polymeric flux $\Phi_{p}^{\delta q^{2}}$ dominates the Lamb vector has a larger likelihood of being normal to the local advecting velocity, with the distribution still being right skewed. This agrees with the fact that in this range of scales, only a small fraction of energy is transferred to smaller scales by the inertia-driven transfer route. The weakening of the $\text{d}\Phi_{f}^{\delta q^{2}}/\text{d}r$ interscale-energy transfer in the elastic range of scales is thus accompanied by an increase of the events where $\bm{u}^{*}(r)$ and $\delta\bm{\omega}(r)\times\delta\bm{u}(r)$ are perpendicular. When considering large $r$, where the influence of the polymers is weaker and $\Phi_{f}^{\delta q^{2}}$ dominates, the distribution of $\cos(\theta)$ flattens and progressively recovers a distribution similar to that in HIT.
5 Scale enstrophy
5.1 Scale-by-scale budget
Figure 9: Scale-by-scale enstrophy budget. In order from left to right and from top to bottom, the panels are for $De=0$, $De=1/9$, $De=1/3$, $De=1$, $De=3$ and $De=9$.
We plot the scale-by-scale budget for $\left\langle{\delta\omega^{2}}\right\rangle$ in figure 9. Unlike energy and helicity, enstrophy is not an inviscid invariant of the Navier–Stokes equations. The budget equation for $\left\langle{\delta\omega^{2}}\right\rangle$, indeed, features a source term $\langle V_{st}\rangle$, that quantifies the amount of enstrophy that is created up to the scale $r$ by means of vortex stretching-like processes. This means that even in HIT, the notion of cascade used for energy transfer does not apply here, and most of the enstrophy is generated directly at the small scales. However, as stated by several authors (see for example Davidson et al., 2008), vortex stretching does occur in the inertial range of HIT and hence a net transfer of enstrophy to smaller scales is expected at these $r$.
We start looking at HIT ($De=0$). Figure 9 shows that in the inertial range, vortex stretching is balanced by the dissipation of enstrophy, i.e., $\langle V_{st}\rangle(r)\approx(4/3)\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle r$, resulting in the $\langle V_{st}\rangle(r)\sim r^{1}$ scaling. This can be easily seen by rewriting equation 18 for Newtonian turbulence, by dropping the polymeric contributions. In this range of scales the forcing $\langle F^{\delta\omega^{2}}\rangle$ and dissipation $\langle D_{f}^{\delta\omega^{2}}\rangle$ contributions to the total flux $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$ disappear, and $\langle\Phi_{f}^{\delta\omega^{2}}\rangle\ll\left\langle{V_{st}^{\delta\omega^{2}}}\right\rangle$ as shown in figure 9: we indeed have that
$\left\langle{V_{st}^{\delta\omega^{2}}}\right\rangle(r)\approx-\frac{4}{3}\left\langle{\varepsilon_{f}^{\delta\omega^{2}}}\right\rangle r.$ (23)
A more insightful way to understand such a linear scaling behaviour of $\left\langle{V_{st}^{\delta\omega^{2}}}\right\rangle$ is by realising that vortex stretching receives its maximum contribution from the smallest scales. This can be seen by the dimensional estimate of the scale-by-scale vorticity $\omega_{r}$

### v2

@@ 4.1 Scale energy $\left\langle{\delta q^{2}}\right\rangle$
We now examine the influence of polymeric additives on the energy transfers. We begin by analysing the budget equation for $\left\langle{\delta q^{2}}\right\rangle$, focusing on how the average energy cascade is altered. This real-space analysis complements the Fourier-space study conducted by Rosti et al. (2023). Subsequently, we explore the nonlinear and polymeric fluxes in detail, highlighting the role of polymers in modifying the intermittency of energy transfers and the localised backscatter.
4.1.1 Scale-by-scale budget
@@
Figure 6: All contributions in budget equation for $\left\langle{\delta q^{2}}\right\rangle$. The panels are for $De=0$, $De=1/3$, $De=1$, and $De=9$, from top-left to bottom-right. The yellow and blue shaded regions identify the inertial and elastic ranges of scales. Here $\Phi_{t}=\Phi_{f}+\Phi_{p}$ and $D_{t}=D_{f}+D_{p}$.
Figure 6 represents the terms in the budget equation for $\left\langle{\delta q^{2}}\right\rangle$ for both Newtonian HIT and PHIT.
@@
We begin by examining the $De=0$ case, which represents the classical Richardson-Kolmogorov turbulence scenario for Newtonian fluids. Energy is injected at the largest scales, $r>r_{IE}\approx 1.56$, through external forcing, with $\langle F^{\delta q^{2}}\rangle\approx(4/3)\langle\varepsilon_{f}^{\delta q^{2}}\rangle r$. From there, energy is transferred from larger to smaller scales within the inertial range of scales $0.028\approx r_{DI}\leq r\leq r_{IE}$ via the classical inertia-driven cascade. Here, $r_{IE}$ marks the scale above which forcing dominates, separating the inertial and energetic ranges, while $r_{DI}$ denotes the scale below which viscous dissipation dominates, separating the dissipative and inertial ranges. Consistent with the Kármán-Howarth equation, the energy transfer rate remains constant within the inertial range and equal in magnitude to the dissipation rate as evidenced by the plateau in the top-left panel of figure 6, here $\langle\Phi_{f}^{\delta q^{2}}\rangle\approx-(4/3)\langle\varepsilon_{f}\rangle r$,. At scales smaller than $r_{DI}$, viscous effects $\langle D_{f}^{\delta q^{2}}\rangle$ dominate, and $\langle\delta q^{2}\rangle$ is dissipated into heat through viscous friction.
@@
Polymeric additives introduce an additional dissipative process $D_{p}^{\delta q^{2}}$ and an alternative route for energy transfer among scales $\Phi_{p}^{\delta q^{2}}$. For small $De\ll 1$, the global picture resembles what observed in Newtonian HIT. In this case, indeed, the polymers are only marginally stretched as their relaxation time is small compared to the characteristic time scales of the flow, and they quickly relax back to their equilibrium lengths; here, $\Phi_{p}^{\delta q^{2}}\approx D_{p}^{\delta q^{2}}\approx\pi^{\delta q^{2}}\approx 0$.
@@
As $De$ increases towards $\mathcal{O}(1)$ the relevance of the non-Newtonian contribution increases, and the polymeric additives modify the energy cascade in a non trivial way. Energy enters the system at the large scales ($r>r_{IE}$) and is dissipated away at the same rate, via both the fluid dissipation and the polymeric sink term, i.e. $\langle F^{\delta q^{2}}\rangle=(4/3)(\langle\varepsilon_{f}^{\delta q^{2}}\rangle+\langle\pi^{\delta q^{2}}\rangle)r$. In the intermediate range of scales ($r_{DI}<r<r_{IE}$) energy is then transferred from larger to smaller scales by two alternative routes, respectively associated with (i) the classic inertia-driven term $\Phi_{f}^{\delta q^{2}}$ and (ii) the fluid-polymer interaction term $\Phi_{p}^{\delta q^{2}}$. At these scales the viscous effects are negligible, and the total energy transfer rate is an invariant and matches the total dissipation rate (Chiarini et al., 2025); here $\langle\Phi_{f}^{\delta q^{2}}\rangle+\langle\Phi_{f}^{\delta q^{2}}\rangle\approx-(4/3)(\langle\varepsilon_{f}^{\delta q^{2}}\rangle+\langle\pi^{\delta q^{2}}\rangle)r$. Eventually, at the smallest scales ($r<r_{DI}$) energy is dissipated away by both the fluid and the polymers as $\text{d}\langle D_{f}^{\delta q^{2}}\rangle/\text{d}r\rightarrow(4/3)\langle\varepsilon_{f}^{\delta q^{2}}\rangle$ and $\text{d}\langle D_{p}^{\delta q^{2}}\rangle/\text{d}r\rightarrow(4/3)\langle\pi^{\delta q^{2}}\rangle$ for $r\rightarrow 0$.
@@
The range of scales over which the energy cascade is significantely altered by the presence of polymers varies with $De$, as indicated by the blue shaded region in figure 6. To quantify this, we define for each $De$ a characteristic scale $r_{p}^{*}$ below which elastic effects are expected to dominate. This scale is estimated by comparing the local eddy turnover time at scale $r$, i.e. $\tau_{f}(r)\sim r/\left\langle{\delta V^{2}}\right\rangle^{1/2}$, where $\delta V(r)=(u_{i}(\bm{x}+\bm{r})-u_{i}(\bm{x}))\cdot r_{i}/r)$, with the polymeric relaxation time $\tau_{p}$. Thus, $r_{p}^{*}$ is estimated as
@@
For $r>r_{p}^{*}$ the time scale of the fluid fluctuations is larger than $\tau_{p}$ ($\tau_{p}/\tau_{f}(r)<1$): the polymers are only marginally stretched and thus their influence on the flow is weak. On the contrary, when $r<r_{p}^{*}$ the time scale of the fluid is smaller than the polymeric relaxation time ($\tau_{p}/\tau_{f}(r)>1$): the polymers interact more effectively with the fluid and thus modify the organisation of the corresponding fluctuations. For the present cases, we measure $r_{p}^{*}\approx 0.04,0.32,1.34$ and $2.81$ for $De=1/9,1/3,1$ and $3$: the range of scales where elastic effects significantly modify the energy cascade increases with $De$, in agreement with the widening of the elastic range of scales shown in figure 6. For $De=9$, $\tau_{p}>r/\left\langle{\delta V^{2}}\right\rangle^{1/2}$ for all $r$, suggesting that the polymeric additives effectively modify the energy cascade in the whole range of scales. The scaling behavior observed in figure 2 aligns with $r_{p}^{*}$ lying well within the inertial range $r_{DI}<r\leq r_{IE}$ when $De$ and $Re$ are sufficiently large. For small $De$ and/or low $Re$, the $\langle\delta q^{2}\rangle\sim r^{1.3}$ scaling disappears as $r_{p}^{*}$ falls within the dissipative range (De Angelis et al., 2005; Perlekar et al., 2010). We reiterate that when $De$ is very large and $\tau_{p}\gg r/\left\langle{\delta V^{2}}\right\rangle^{1/2}$ the polymeric chains decouple from the carrier phase and their influence on the carrier flow is rather low.
This suggests that the relevant control parameter is not the global Deborah number, but rather the scale-dependent or local Deborah number, defined as $De(r)=\tau_{p}/\tau_{f}(r)$, where $\tau_{f}(r)$ is indeed the characteristic timescale of eddies of size $r$. This perspective is consistent with the observation that different bulk flow quantities exhibit a change in behaviour at different values of (the global) $De$; see table 1. Specifically, the inversion in the behaviour of a given quantity due to the change of regime is expected to occur when $De(r)$ becomes sufficiently large at the scale $r$ where that quantity is most active or sensitive.
A last comment regards the influence of $De$ on the dissipative range of scales. Our data show that $r_{DI}$ (i.e. scale below which the total dissipative term $D_{t}^{\delta q^{2}}=D_{f}^{\delta q^{2}}+D_{p}^{\delta q^{2}}$ dominates) decreases with $De$, thus indicating that the dissipative range of scales progressively shrinks as the polymeric relaxation time increases.
4.1.2 The nonlinear and non Newtonian fluxes
Refer to captionRefer to caption
@@ 4.1.2 The nonlinear and non Newtonian fluxes
Refer to captionRefer to caption
@@
Figure 8: Dependence of the distribution of the non Newtonian flux $\Phi_{p}^{\delta q^{2}}$ on the Deborah number, for (left) $r=0.1$, (centre) $r=0.3$, and (right) $r=0.9$. Top: distribution of $\Phi_{p}^{\delta q^{2}}$. Bottom: distribution of $\Phi_{p}^{\delta q^{2}}/\Phi_{p,rms}^{\delta q^{2}}$. In the bottom panels the inset provide a zoom of the normalised distributions.
@@
The average picture of the cascade and interscale exchanges in the $r$-space is not representative of the actual physical processes. In the previous section, we have shown the influence of the polymers on the energy cascade in an average sense only, even though it is known that the interscale energy transfer is highly intermittent (see for example Piomelli et al., 1991; Domaradzki et al., 1993; Cerutti & Meneveau, 1998). Moreover, even though energy cascade in HIT is on an average from larger to smaller scales, there exist localised regions in the flow where energy actually cascades from smaller to larger scales, opposite to the average sense. We now investigate how the presence of polymeric additives modifies this picture, and whether this localised inverse energy transfer is also detected in the transfer mechanism driven by the fluid-polymer interaction.
@@
Figures 7 and 8 show the probability distribution functions of the nonlinear energy fluxes $\Phi_{f}^{\delta q^{2}}$ and the polymeric energy fluxes $\Phi_{p}^{\delta q^{2}}$ for various values of $De$. We analyse three different separations: $r=0.1$, where $\Phi_{p}^{\delta q^{2}}$ dominates for $De>1$, and $r=0.3$ and $r=0.9$ where $\Phi_{f}^{\delta q^{2}}$ dominates for all $De$. In our convention, a positive flux $\Phi>0$ corresponds to backscatter events, i.e., local transfers of energy from smaller to larger scales, whereas a negative flux $\Phi<0$ indicates forward cascade events, transferring energy from larger to smaller scales. The distributions reveal the highly intermittent and strongly non-Gaussian nature of both $\Phi_{f}^{\delta q^{2}}$ and $\Phi_{p}^{\delta q^{2}}$, being characterised by heavy tails (Ishihara et al., 2009; Yasuda & Vassilicos, 2018). Both extreme forward transfers and backscatter events occur with significantly higher probability than predicted by a normal distribution. The distributions are skewed towards negative values, confirming that on average both fluxes transfer energy forward across scales. Quantitatively, for $De=1$ at $r=0.9$ ($r=0.1$) the probability that $\Phi_{f}^{\delta q^{2}}$ exceeds its root-mean-square value is $0.0672$ ($0.0274$), while the probability that it falls below minus its root-mean-square value is $0.1640$ ($0.0767$). Similarly, for $\Phi_{p}^{\delta q^{2}}$, the probabilities of exceeding and being less than minus its root-mean-suqare value are $0.0323$ ($0.0105$) and $0.2402$ ($0.1545$), respectively.
We now examine how polymers affect the tails of the distributions, focusing on extreme events. Starting with the nonlinear flux $\Phi_{f}^{\delta q^{2}}$ shown in figure 7, we observe that across all three scales, the presence of polymers narrows the tails of the distribution. This indicates a suppression of extreme forward and backward energy transfer events along the classical cascade route. This attenuation becomes more pronounced as $De$ increases, consistent with a smoother velocity field and the diversion of energy towards the polymer-driven transfer mechanism. When examining the fluxes normalised by their root-mean-square values (bottom panels), we find that the distributions at $r=0.3$ and $r=0.9$ collapse reasonably well across all $De$. Some deviations appear at the largest values, likely due to insufficient statistical convergence. This suggests that, in the $\Phi_{f}^{\delta q^{2}}$-dominated range, polymers modify the intensity of extreme nonlinear cascade events while having little effect on their intermittency. In other words, in the inertial range of scales where the polymeric contribution is weak, the self-similarity of the classical cascade process appears to be largely maintained. On the contrary, the normalised distributions do not overlap at the small scales $r=0.1$. In the elastic range of scales the nonlinear flux $\Phi_{f}^{\delta q^{2}}$ has a small average contribution, and its intermittency increases with $De$. We now move to the polymeric flux $\Phi_{p}^{\delta q^{2}}$, shown in figure 8. As $De$ increases, the distribution tails widen, and the asymmetry becomes more pronounced consistent with the rise in the average value. When examining the normalised fluxes (bottom panels), the behaviour differs from that of $\Phi_{f}^{\delta q^{2}}$. Specifically, the distributions collapse reasonably well at the small scale $r=0.1$ (with some deviations at extreme values), but not at the larger scales $r=0.3$ and $r=0.9$. An increase in $De$ appears to alter the space-time intermittency of the interscale energy transfer driven by the polymeric microstructure primarily at larger scales, while self-similarity with respect to $De$ appears to hold at smaller scales within the elastic range of scales. Consistent with the increased asymmetry, at large scales higher $De$ corresponds to narrower positive tails and slightly heavier negative tails.
[…]
@@
We now provide a further insight on the influence of the polymers on the inertia-driven energy cascade process by looking at the simple relation introduced by Baj et al. (2022), that links the nonlinear flux with the Lamb vector $\delta\bm{\omega}\times\delta\bm{u}$. Exploiting the Lamb decomposition and manipulating the equations, it is indeed possible to show that
@@
[…]